### v2

# Tracking electron capture process in classical molecular dynamics simulations for spectral line broadening in plasmas

Abstract
Plasma spectroscopy is a fundamental tool for diagnosing laboratory and astrophysical plasmas. Accurate interpretation of spectra depends upon precise modeling and comprehension of Stark-broadening and other mechanisms affecting spectral lines. In this context, computer simulations have emerged as valuable tools, offering idealized experiments with well-defined conditions. Molecular dynamics simulations, in particular, excel at replicating the particle interactions within the plasma and their impact on the state of a radiating atom or ion. However, these simulations present challenges in tracking electron capture processes, since setting an unambiguous criterion to distinguish between bound and free electrons is not trivial. In this paper we introduce a new algorithm that, within a classical framework, precisely identifies the scenario in which an electron is captured by an ion and then follows a stable orbit around it. The algorithm’s applicability extends to emitters with charges $Z\geq 1$. The procedure enables the correct identification of valid time-histories of the electric microfield perturbing the emitting ion, which will be used for subsequent line shape calculations. The ionization balance results obtained from the application of this algorithm are compared with an additional method based on the potential energy of the particles in the simulations, finding good agreement, therefore validating the use of this approach.
I Introduction
Accurate modeling of spectral line shapes is indispensable for investigating the properties of high energy density plasmas. The analysis of emission or absorption spectra, which typically involves the comparison of experimental data with spectroscopic models, stands as a key diagnosis tool for laboratory and astrophysical plasmas. In particular, spectral line shapes, broadened and shifted by the kinetics of plasma particles and their interactions, provide relevant information about the plasma temperature and density. Spectral lines arise from atomic transitions between bound energy states of a radiating ion or neutral atom, the radiator.
In the context of hot and dense plasmas the Stark effect emerges as the dominant broadening mechanism [1]. Unlike the isolated scenario, when embedded in a plasma, the radiator is subject to the electric field generated by its surrounding charged particles during the time interval in which it emits. The shape of the spectral lines in this case depends on the plasma-induced electric field, its time evolution (tied to the plasma particle dynamics) and the radiators’ ensemble.
Despite significant strides in recent decades [2, 3], modeling spectral line shapes remains a formidable challenge owing to its interdisciplinary nature, requiring the integration of various ingredients from atomic physics, plasma physics, statistical mechanics, collision theory, and computational physics. Moreover, in the high-density regime, dedicated validating experiments are notably scarce, presenting a substantial obstacle to grounding theoretical models in empirical reality.
Generally, two primary approaches are used for line-broadening calculations: analytic methods and computer simulations. Analytic models navigate the complexities of line broadening by incorporating assumptions and mathematical manipulations to handle the stochastic nature of line emission. One common approximation involves the separate calculation of electron and ion broadening contributions. Since ions are much heavier than electrons, they are assumed to be static over the relevant electron time scales. The analytic approach then puts the effort in obtaining an accurate representation of the electron broadening operator. In this regard, some common approximations include the simple and well-known impact approximation, or the second-order approximation as a simplified way to compute the electron broadening operator [1, 2]. Once the electron broadening contribution is computed for a given ion distribution, i.e. for a given value of the ion microfield, the line emission is integrated over the statistical distribution of the ion microfield for the conditions of interest, a facet that entails its own extensive research area in microfield probability distributions. The inherent advantages of analytic methods lie in their computational efficiency, as the time required to calculate a line shape is generally much shorter than for computer simulations, and they operate without numerical noise.
On the other hand, computer simulations prove particularly advantageous in gaining insights into the line broadening problem, excelling at replicating plasma particle interactions and their impact on the state of a radiating atom or ion [4]. Simulated plasma particles traverse within a confined space, following classical trajectories in discrete timesteps. These classical particles generate a time-dependent Coulomb or Debye potential or, equivalently, a time-sequence of the electric field. The emitting atom or ion then internally evolves subjected to the effect of this electric field. By numerically solving the time-dependent Schrödinger equation, one obtains the corresponding evolution operator. It is then possible to calculate the dipolar moment of said emitter as a function of time. The Fourier transform of the autocorrelation function of this dipolar moment corresponds to the emission spectrum from the emitter under consideration. By repeating this process a sufficient number of times, the spectrum is averaged over an ensemble of emitters producing the spectral line of interest.
Simulations often simplify the problem and reduce computational time by employing straight-line path trajectories for particle movement, i.e. with the simulated particles being independent. We will refer to this type of simulations as Trivial Molecular Dynamics (TMD). TMD calculations are useful when the typical electric-interaction energies are much lower than the mean kinetic energy of particles (a weakly-coupled high temperature plasma) and, specially for non-charged emitters [5, 6, 7, 8, 9, 10].
Noteworthy exceptions to this norm are Full Molecular Dynamics (FMD), in which the calculation takes into account the electric interaction of all particles in the plasma. This constitutes a more accurate (and more computationally expensive) method, which must solve Newton’s equations of motion to determine the trajectories of all particles. FMD simulations are necessary for medium and strongly-coupled plasmas and, specially when the emitter is an ion, given that its interaction with the surrounding particles in the plasma affects both its own trajectory and those of the other particles, thus modifying the time-evolution of the electric field of interest [11, 12, 13]  [^1] It is worth mentioning that there is an intermediate type of calculations, in which the perturbers are treated as independent particles, but their interaction with the emitter is taken into account explicitly. This results in the charged perturbers moving in hyperbolic trajectories when a Coulomb field is considered, or in slightly different paths if Debye screening is introduced, in which case the corresponding differential equation must be solved numerically [14]. .
While computer simulations have become a standard tool to study Stark-broadened lineshapes, there are some Physical broadening mechanisms that are still unaccounted for in these type of calculations. In this work we focus on the so-called recombination broadening, produced when a free electron within the plasma becomes bound to an emitting ion, therefore stopping its emission. While it is clear that TMD models cannot reproduce this phenomenon, as the particles are considered independent, it is possible to use FMD simulations to detect when an electron becomes bound to an ion and account for this additional broadening.
It is in this context that our work is framed. We present an algorithm to accurately track the electron capture processes within FMD computer simulations, key to obtain the correct time-sequences of the electric field within the plasma and accurately account for the additional recombination broadening.
This manuscript is structured as follows: in Section II we present the criterion for detecting electron captures within FMD simulations. An example of the application of this criterion to a computer simulation is presented in Sec. III.1, and the results are compared and validated with an additional metric in Sec. III.2. Finally, Sec. IV summarizes the conclusions and limitations of this work and explores future applications of this type of techniques.
II Criterion for detecting trapped electrons in a computer simulation
As mentioned in Section I, in FMD simulations, the particle trajectories are solved by accounting for the Coulombian interaction between all particles in the plasma. As the simulation evolves, it becomes possible that the binding energy (negative) between an electron and an ion becomes too large - in absolute value - with respect to their kinetic energies (positive), so that the electron becomes trapped by the ion. The two particles in this pair can therefore no longer be considered free and must be labeled as a neutral atom, or as an ion with a positive charge lower than that of the free ion.
This circumstance was extensively discussed in our previous work [15], where we developed a statistical model for the equilibrium of the simulated plasma, taking into account the process of electron capture and release by the ions (recombination and ionization respectively). For this model, it was necessary to establish a clear criterion to determine when an electron-ion pair must be considered bound, and when those particles must be considered free. While in Gigosos et al. [15] ions of charge +e+e were used (where $e$ is the electron charge), in this work we are interested in plasmas with charge +Ze+Ze, with $Z>1$. The purpose of this work is then to extend this criterion to situations in which the number of trapped electrons can exceed $1$.
The need for such a criterion is obvious. In the simulations, it is necessary that the density of free electrons is a given value (generally predetermined). To calculate such value, one needs to count the electrons that are really free in the simulated plasma. This number does not match that of the electrons included in the calculations, since some of them will become bound to the ions.
Furthermore, there is another important reason that requires knowing when an ion has captured an electron and when it can be considered free. Let us consider a FMD simulation in which we are interested in calculating the spectrum from an ionic emitter with charge +Ze+Ze. To that end, during the simulation process, we must register the temporal sequence of the electric field seen by that charged emitter, which is produced by the ions and free electrons surrounding it. If, at a given moment, an electron is captured by the ion of consideration, from that moment on, the electron will orbit around it at a short distance. Therefore, the electric field felt by that emitter will be an oscillating function with a large amplitude (since the oscillating electron is very close), which spoils the field sequence.
In addition, the capturing ion, originally with charge +Ze+Ze becomes now, along with the trapped electron, an ion of charge +(Z1)e+(Z-1)e and can therefore no longer be considered a contributor to the spectral line in which we are interested. The moment the electron is captured determines when the ion changes its species and therefore the emission coherence is abruptly lost. This causes a broadening of the line by the recombination process, which highlights one the advantages of these calculations [13]. As a consequence, in practice, the corresponding time-sequence of the electric field is cut in the simulation process. It is then necessary to wait for the corresponding ion to recover its original charge in order to continue using it for the calculation.
Usually, electrons are labeled as bound to an ion when their distance is lower than a given value $r_{min}$ (whose value is discussed in Gigosos et al. [15]) and the total energy (kinetic plus potential) of the electron-ion pair is negative. While this criterion [16] works well in most cases, it gives some wrong answers. Moreover, for ions with charge $Z>1$, this method only detects if an ion has some electron trapped, but not how many [17].
If one follows this criterion, it is possible for an electron to orbit at distances larger than $r_{min}$ (and thus be considered free), but still produce a non-physical oscillating electric field on the ion. On the other hand, for strong enough collisions (an electron gets closer than $r_{min}$ to an ion, but only momentarily), the total energy of the electron-ion pair can become negative, which, following this criterion, would be wrongly considered as them being bound. Since these collisions happen regularly, this criterion effectively eliminates long field sequences of the Stark calculation, as they are broken every time a false bound pair is detected. Furthermore, under this criterion, the field sequences considered will underestimate the effect of strong collisions. These two effects become more important for ions with charge $Z>1$ (as more electrons can become bound), and more strongly coupled plasmas.
To correctly select the valid field sequences we choose a global criterion, in which whether an ion has trapped an electron or not, not only depends on its conditions at a given time, but also on its history. For a given ion, we store (for each step of the simulation) its potential energy, the electric field at its position and the label (an index identifying each electron in the simulation) of the $Z+1$ closest electrons, where $Ze$ is the charge of the emitter.
In order to correctly calculate the potential energy, since the interaction between particles is Coulombian, it is necessary to apply a regularization to the interaction potential to avoid divergencies at short distances. While several procedures for this regularization have been reported (see Bonitz et al. [18] and references therein), in this work we will use that described in Gigosos et al. [15], which considers a parabolic potential for interparticle distances below a given value $a$ and hyperbolic for larger distances. This method naturally accounts for the diffraction of electrons around an ion and determines an effective atom size. We used the following expression for $Z>1$ ions, rescaled from Gigosos et al. [15]
$V_{ie}(r)=\begin{cases}V_{i}\left[\frac{1}{3}\left(\frac{r}{a}\right)^{2}-1\right],&\quad r\leq a\\ -\frac{Ze^{2}}{4\pi\varepsilon_{0}r},&\quad a<r\leq R_{I}\\ 0,&\quad r>R_{I}\end{cases}$ (1)
where $V_{i}$ is the ionization energy corresponding to the ion of charge $Z$ and $R_{I}$ is the radius of the interaction sphere, defined as half the size of the cubic simulation box. For ions with $Z>1$, $a$ is a parameter given by
$a=\frac{3}{2}\cdot\frac{Ze^{2}}{4\pi\varepsilon_{0}V_{i}}.$ (2)
The first step then is to track down the moments when the potential energy of the emitter is lower than a given threshold energy $\mathscr{E}_{th}$. In particular, to keep consistency with the original criterion for $Z=1$ atoms, we choose
$\mathscr{E}_{th}=-\frac{2}{3}V_{i},$ (3)
which corresponds to the potential energy of an ion with an electron at a distance $a$. We note here that the exact value of $\mathscr{E}_{th}$, while important, is not critical for this criterion, as we will show in Section III.1.
Whenever the ion energy is lower than $\mathscr{E}_{th}$, the electron that is closest to it (its first neighbor) is identified. Then, the time interval during which that electron has been among the $Z+1$ closest neighbors is considered. If this interval is longer than a given threshold value $\tau_{\text{bound}}$, then the average total energy of the electron-ion pair in that interval is calculated. If the pair has a negative total energy during that interval, then the electron is considered to be bound for the whole interval. If not, it is considered that the negative energy seen by the emitter is caused by a strong collision, since the electron causing it did not remain around the ion. It is worth noting that the $Z+1$ closest neighbors are considered, since it is possible that a fast electron gets closer to the ion core than other bound electrons for a short period of time, without freeing them.
Since the full time-history of the ion is needed in order to check whether an electron is bound or free, its history must be saved so that it can be accessed when the energy criterion is met, and the correct field sequences can be extracted.
Determining the timescale for bound electrons
In the classical picture of these FMD simulations, an electron orbiting an ion will do so with a given period, such that the orbit is stable. For a hyperbolic potential (such as the Coulomb potential), the condition for a stable orbit is given by
$m_{e}v^{2}=\frac{Ze^{2}}{4\pi\varepsilon_{0}r},$ (4)
where $m_{e}$ is the electron mass, $v$ its velocity, and $r$ the orbit radius. Under this condition, an electron orbiting an ion at the typical thermal velocity ($v_{T}=\sqrt{2k_{B}T_{e}/m_{e}}$, with $k_{B}$ the Boltzmann constant and $T_{e}$ the electron temperature) will do so at a distance of
$r_{T}=\frac{Ze^{2}}{8\pi\varepsilon_{0}k_{B}T_{e}},$ (5)
and therefore, the corresponding orbital period is given by
$\tau_{T}=\frac{2\pi r_{T}}{v_{T}}=\frac{Ze^{2}}{2\varepsilon_{0}}\cdot\frac{m_{e}^{1/2}}{(2k_{B}T_{e})^{3/2}}.$ (6)
This expression can be simplified by defining the typical timescale of electron dynamics $t_{0}$ as the time it takes an electron moving with the characteristic thermal velocity $v_{T}$ to travel the typical distance between electrons ($r_{e}=(4\pi n_{e}/3)^{-1/3}$). Using this definition, one gets
$t_{0}=\frac{r_{e}}{v_{T}}=\left(\frac{3}{4\pi n_{e}}\right)^{1/3}\cdot\left(\frac{m_{e}}{2k_{B}T_{e}}\right)^{1/2},$ (7)
which yields
$\tau_{T}=\frac{\pi}{3}Z\rho^{2}t_{0},$ (8)
where $\rho$ is the coupling degree of the plasma, given by the relation between the typical distance between electrons and the Debye radius.
We take this characteristic orbit time as a basis to consider whether an electron is bound to an ion. Although one might be tempted to consider $\tau_{T}$ as the temporal threshold - that is, if an electron is within the $Z+1$ closest neighbors of a given ion for at least $\tau_{T}$, and its total energy is negative, it is considered bound - this criterion would overestimate the number of bound electrons, since an electron can loop around an ion without being bound. To account for this possibility, we take $\tau_{\text{bound}}=3\tau_{T}$ as the threshold for considering an electron as bound to a given ion.
We note here that the above expression for $\tau_{T}$ was obtained for a hyperbolic potential. However, as indicated in Eqn. 1, in our simulations, the potential becomes parabolic for distances $r<a$, and thus, for this expression to be valid, it is necessary that $r_{T}>a$. By comparing the above expressions it can easily be seen that this condition is equivalent to
$\frac{V_{i}}{2}>\frac{3}{2}k_{B}T_{e}.$ (9)
Therefore, the criterion is valid if the ionization potential is larger than twice the thermal energy of electrons. However, this is always the case when recombination effects are relevant given that, in reality, if the temperature is high enough so that it becomes comparable with the ionization potential, the electrons are all free, and recombination effects are negligible (as we will show in the following section).
Figure 1: Time-history of a plasma with $\rho=0.6$ and emitters with charge 2. From top to bottom: potential energy of an emitter as a function of time, indices of the three closest electrons at each moment, and electric field. The shaded regions correspond to the moments when a particle is trapped. The potential energy is given in units of $\mathscr{E}_{0}$, a characteristic energy of the simulation, and the electric field in units of $E_{0}$, the electric field produced by a single electron at the typical distance $r_{e}$.
III Results and discussion
III.1 Application of the criterion
Let us now show an example of the application of this method, for the particular case of ions with charge $Z=2$, which means that the three closest neighbors must be monitored. In this example, the coupling parameter of the plasma is $\rho=0.6$, and therefore $\tau_{T}\sim 0.75t_{0}$ and $\tau_{\text{bound}}\sim 2.25t_{0}$. Figure 1 shows the time-history of the potential energy of an emitter (top), the indices $I_{p}$ of the three electrons that are closest to it at every timestep in the simulation (middle) and the value of the electric field acting on said emitter (bottom panel). The potential energy is given in units of the characteristic energy of the simulation $\mathscr{E}_{0}$, a parameter given by the average kinetic energy of the particles at the beginning of the simulation –before equilibrium is reached–, as further detailed in Gigosos et al. [15]. The electric field is given in units of $E_{0}$, i.e. the field produced by a single electron at the typical distance $r_{e}$.
At $t\sim 2.7t_{0}$, when the potential energy becomes lower than $\mathscr{E}_{th}$, the electron that is closest to the emitting ion is electron number 298. We then observe that this electron has been among the three closest neighbors from $t\sim 0$ to $t\sim 10t_{0}$, approximately 10 times the characteristic time of the simulation. This is much longer than the typical duration of a collision, so we must assume that this corresponds to a bound electron. Indeed, if we now consider the electric field between $t\sim 0$ and $t\sim 10t_{0}$, it oscillates periodically, with a period of approximately $0.75t_{0}$ (which is the value of $\tau_{T}$ in this case), reaching very high values. A similar situation is observed at $t\sim 40t_{0}$, although in this case the trapped electron oscillates around the ion only a few times, and therefore its imprint on the electric field history is not as clear.
It is clear now that the exact value of $\mathscr{E}_{th}$ is not critical, since strong collisions and electron captures appear in the time-history of the potential energy of the ion as clear spikes, and what determines whether an ion has captured an electron or not, is the time-dependent part of the criterion. In the example presented in Fig. 1, the ionization potential is $V_{i}=15\mathscr{E}_{0}$, which following the definition in Eqn. 3 corresponds to $\mathscr{E}_{th}=-10\mathscr{E}_{0}$. However, if its value were taken between $\sim-3\mathscr{E}_{0}$ and $\sim-12\mathscr{E}_{0}$, the same results would have bee obtained.
In fact, by taking a lower threshold energy, the only difference is the computational cost of the simulation, since the time-history of the ion would need to be checked more often. For example, at $t\sim 20t_{0}$, the potential energy jumps below $6\mathscr{E}_{th}$, as an electron passes very close to the emitting ion. If the chosen threshold energy were lower, the algorithm would inspect this event as a potential electron capture. However, the colliding electron in this case does not remain among the three closest neighbors for long enough to be a bound particle. Indeed, the electric field history confirms that this corresponds to an isolated collision. It can therefore be seen that this system does not eliminate strong collisions and thus, by applying this criterion, it is possible to identify and select the appropriate sequences of the electric field that a radiator is subjected to during its emission, including the correlation loss owed to capturing an electron. These sequences can then be further used for Stark-broadening line shape calculations.
III.2 Validation of the trapped-electron detection algorithm: calculation of the ionization balance
Figure 2: Distribution of the potential energy of ions (left) and its integral (right). Data correspond to a Helium plasma ($Z=2$), with an ionization potential $V_{i}=30\,\mathrm{\mathscr{E}_{0}}$, and a coupling parameter $\rho_{N}=1.0$. The mean total energy per particle of the system is $\mathscr{E}_{t}=-2.12\mathscr{E}_{0}$.
Figure 3: Same as in Fig. 2, with a mean total energy for particle of $10.00\mathscr{E}_{0}$.
In order to validate the electron-trapping detection algorithm that we have established, we aim to compare its predictions with those from a different independent method, consistent with the FMD framework. In this regard, we note that this algorithm can additionally be used to determine the number of electrons that are trapped by an ion. As such, using this algorithm, one can easily determine the fractional population of different ion charge states in the plasma at given conditions and therefore obtain the corresponding ionization balance.
Nevertheless, the plasma ionization balance can also be obtained from the analysis of the ion potential energy distribution. In a FMD simulation with a given set of conditions, one can easily obtain the potential energy distribution of the plasma particles at any given time. By averaging over a sufficient number of instants and a sufficient number of simulation runs starting with different initial conditions ($\sim 10$), it is possible to obtain a good statistical representation of this distribution.
To illustrate how the ionization balance can be obtained from the potential energy distribution, let us first assume the simple case of hydrogen ions ($Z=1$). If the plasma temperature is low enough so that the distribution of kinetic energies is relatively narrow, the distribution of potential energies presents two main lobes. One of them centered around zero (free particles), and one of them near the potential well of energy $V_{i}$ (ions with a bound electron) – as shown in figure 14 in Gigosos et al. [15] –. In the same manner, we could expect that, for ions with $Z=2$, the distribution of potential energies can have up to three lobes: that corresponding to free particles, around zero potential energy; the one accounting for ions with one bound electron, with potential energies similar to $V_{i}$; and the lobe corresponding to ions with two trapped electrons, which will have energies similar to $2V_{i}$ (and equivalently for higher $Z$).
If these lobes are sufficiently separated, it is possible to integrate the energy distribution and associate the area under each lobe to the population of the different species. This is a robust method that does not require the definition of ad-hoc criteria to label an electron as bound or free. On the contrary, due to its own statistical nature, this method is not useful for selection of electric field sequences since the time-resolved information is lost. Hence, the comparison of the ionization balance obtained from the trapped-electron detection algorithm against the one from the analysis of the potential energy distribution provides a way to validate the former, adding confidence on the selected electric field sequences for line shape calculations.
This is illustrated in Figure 2, in which we show the distribution of potential energies (left) and its integral (right) for an equilibrated plasma of $Z=2$. The corresponding ionization potential is $V_{i}=-30\mathscr{E}_{0}$, and the mean total energy per particle is $2.12\mathscr{E}_{0}$. In this case, the energy per particle is relatively high compared with the ionization potential, and therefore only two ionization species are present in the plasma: free ions and those with one bound electron. The distribution of potential energies therefore only presents two lobes, centered around zero and $V_{i}$. By looking at the integral of the energy distribution (plot on the right), the ratio of ions with one bound electron ($I_{1}$) to free particles ($I_{0}$) can be easily obtained.
A similar example is shown in Figure 3 for the case when the plasma mean energy is $10\mathscr{E}_{0}$. The rest of the plasma parameters are identical as those in Fig. 2. In this case, the plasma energy is too low, and the plasma is composed of only ions with one ($I_{1})$ or two bound electrons ($I_{2}$), with almost no free ions.
Figure 4: Fractional population of the different ion species as a function of the temperature for a plasma of ions with charge 2, and an ionization potential of $V_{i}=30\mathscr{E}_{0}$. Each dot corresponds to a set of simulations with the same temperature at equilibrium.
Using this method, we calculated the fractional population distribution for a collection of simulation runs spanning over a wide range of plasma temperatures for the case of $Z=2$ and $V_{i}=-30\mathscr{E}_{0}$. The results are shown in Figure 4, where the the different colors correspond to the population fraction of the different ion species as a function of temperature. We note that each dot in fact corresponds to a group of different simulations associated to the same temperature value, i.e. for a given temperature, a number of simulations starting from different initial conditions were launched and then let to thermalize following the method described in our previous work [15]. As expected, at low temperatures all ions have two bound electrons and, as the temperature increases, some electrons become free, leading to an increase in the population of ions with only one bound electron. The number of ions with $Q=+1$ keeps growing until almost no neutral ions remain. If the temperature is increased further, ions can become doubly ionized ($Q=+2$), until there are no more bound electron-ion pairs.
Once the ion population distribution is calculated, one can easily obtain the average ionization, which permits the comparison with the results from the application of our trapped-electron detection algorithm. We show this in Figure 5 where the results from the two methods are shown for the same temperature range. The agreement between the two methods is remarkable, which firmly constitutes a validity proof of the electron capture detection algorithm. Additionally, as mentioned before, when the temperature becomes even one tenth of the ionization potential ($\sim 3\mathscr{E}_{0}$), it can be seen that most atoms are fully ionized, and recombination effects become negligible, which validates our estimation of $\tau_{bound}$.
IV Conclusions
Figure 5: Mean ion charge obtained by applying the new criterion for selecting electric field sequences, compared to the values obtained by integrating the potential energy distributions. It can be seen that the results are in very good agreement.
In this study, we have introduced a robust criterion for tracking and identifying trapped electrons in Full Molecular Dynamics (FMD) computer simulations, focusing on ions with charges $Z>1$. Our presented criterion effectively distinguishes strong collisions within the plasma from recombination effects, providing a clear understanding of the underlying processes, and allowing the selection of appropriate time-sequences of the electric field felt by a radiator for Stark broadening calculations. Additionally, this method offers a pathway to determine the plasma ionization balance under specific conditions.
To validate our detection algorithm, we conducted extensive comparisons across a wide range of plasma temperatures. Specifically, we compared the average ionization values obtained through our criterion with those derived from a well-established method based on the analysis of ion potential energy distribution. While our ionization balance calculations do not claim to precisely solve plasma atomic kinetics, our FMD simulations accurately capture the plasma ionization balance within the framework of the classical ionization-recombination mechanism. It is worth mentioning that this classical approach lacks a distinct quantum-mechanical counterpart.
Our simulations, though comprehensive, do not encompass dissipative radiative processes. Consequently, the classical recombination mechanism incorporated in our simulations cannot fully account for radiative recombination; rather, it represents a combined effect arising from electron capture and likely three-body recombination. Nevertheless, the detection algorithm discussed in this study serves as a valuable tool for tracking the ionization-recombination mechanism within full molecular dynamics simulations involving ions with $Z>1$. This inclusion is pivotal, especially when considering the additional recombination width on spectral line shapes.
Importantly, without a methodology akin to the one presented in this study, such effects would be entirely overlooked, leading to inaccurate calculations of line profiles, particularly in strongly-coupled plasmas. By incorporating our detection algorithm, researchers can achieve more precise and reliable simulations, essential for advancing our understanding of spectral line broadening in complex plasma environments.
It is worth mentioning that the criterion presented in this work is applicable when the potential regularization radius $a$ is smaller than the average distance between particles $r_{e}$. While this is most often the case, for heavily coupled systems such as solid-density plasmas [19], care must be taken before enforcing such criterion. Future work on this topic will include the study of FMD simulations techniques for this type of crystalline systems.
Acknowledgements
This work has been supported by Research Grants No. ENE2015-67581-R/FTN (MINECO/FEDER–UE) from the Spanish Ministry of Economy and Competitiveness and PID2022-137632OB-I00 from the Spanish Ministry of Science and Innovation.
This work has also been carried out within the framework of the EUROfusion consortium, funded by the European Union via the Euratom Research and Training Program (Grant Agreement Nos. 633053 and 101052200—EUROfusion). The views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. The involved teams have operated within the framework of the Enabling Research Projects: Grant number AWP21-ENR-IFE.01.CEA. and AWP24-ENR-IFE.02.CEA-01.